# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 7.2 Optimal semiaffine spaces of nonsingular matrices
It remains to understand the nature of the mappings $X\in\mathbb{D}^{n}\mapsto e(X^{\star}PX)$ when $n\geq 1$ and $P$ ranges over the $e$-nonisotropic matrices of $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$. In the special case $d=1$, they are simply the nonisotropic quadratic forms on $\mathbb{D}^{n}$. Hence, as seen in [18], the optimal affine subspaces of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are classified up to equivalence by lists of congruence classes of nonisotropic quadratic forms. In particular, if $\mathbb{F}=\mathbb{R}$ then there is only one such congruence class of $k$-dimensional forms for each $k\geq 1$, and we obtain that every optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ is equivalent to $(I_{n_{1}}+\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{R}))\vee\cdots\vee(I_{n_{p}}+\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{R}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely every space of this form is an optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$.
@@
The situation is not as clear if $d>1$. But say that $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$, and keep the assumption that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$. Then $d\in\{2,4\}$. Let $P\in\operatorname{\operatorname{M}}_{n}(\mathbb{F})$. Then there is a unique splitting $P=Q+R$ where $Q$ is Hermitian (that is $Q^{\star}=Q$) and $R$ is skew-Hermitian. We can then replace $P+\mathcal{M}_{n}$ with $Q+\mathcal{M}_{n}$, and now $X\mapsto X^{\star}QX$ takes all its values in $\mathbb{F}$, so $\forall X\in\mathbb{D}^{n},\;e(X^{\star}QX)=2\,X^{\star}QX$. Moreover, it is known by polarizing that $X\mapsto X^{\star}QX$ determines $Q$. Hence the equivalence class of $P+\mathcal{M}_{n}$ determines $Q$ up to star-congruence and multiplication with an element of $\mathbb{F}^{\times}$. As a consequence, optimal affine subspaces of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are classified, up to equivalence, by lists $([Q_{1}],\dots,[Q_{p}])$ of classes of nonisotropic Hermitian matrices over $\mathbb{D}$, where each class is for the equivalence relation where $Q$ and $Q^{\prime}$ are equivalent if and only if $Q^{\prime}=\alpha R^{\star}QR$ for some invertible $R$ and some $\alpha\in\mathbb{F}^{\times}$. In particular, if $\mathbb{D}\in\{\mathbb{C},\mathbb{H}\}$ and $\mathbb{F}=\mathbb{R}$, then we recover from the classification of Hermitian forms that every optimal $\mathbb{R}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is equivalent to $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal $\mathbb{R}$-affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ for every partition $n=n_{1}+\cdots+n_{p}$.

### v2

@@ 7.3 Making sense of the classification of optimal spaces
It remains to understand the nature of the mappings $X\in\mathbb{D}^{n}\mapsto e(X^{\star}PX)$ when $n\geq 1$ and $P$ ranges over the $e$-nonisotropic matrices of $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$. In the special case $d=1$, they are simply the nonisotropic quadratic forms on $\mathbb{F}^{n}$. Hence, as seen in [12], the optimal unital $\mathbb{F}$-affine subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are classified, up to equivalence, by lists of congruence classes of nonisotropic quadratic forms. In particular, if $\mathbb{F}=\mathbb{R}$ then there is only one such congruence class of $k$-dimensional nonisotropic quadratic forms for each $k\geq 1$, and we obtain that every optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$ is equivalent to $(I_{n_{1}}+\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{R}))\vee\cdots\vee(I_{n_{p}}+\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{R}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely every space of this form is an optimal affine subspace of invertible matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{R})$.
@@
The situation is not as clear if $d>1$. But say that $\operatorname{\operatorname{char}}(\mathbb{F})\neq 2$, and keep the assumption that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$. Then $d\in\{2,4\}$. Let $P\in\operatorname{\operatorname{M}}_{n}(\mathbb{F})$. Then there is a unique splitting $P=Q+R$ where $Q$ is Hermitian (that is $Q^{\star}=Q$) and $R$ is skew-Hermitian. We can then replace $P+\mathcal{M}_{n}$ with $Q+\mathcal{M}_{n}$, and now $X\mapsto X^{\star}QX$ takes all its values in $\mathbb{F}$, so $\forall X\in\mathbb{D}^{n},\;e(X^{\star}QX)=2\,X^{\star}QX$. Moreover, it is known by polarizing that $X\mapsto X^{\star}QX$ determines $Q$. Hence the equivalence class of $P+\mathcal{M}_{n}$ determines $Q$ up to star-congruence and multiplication with an element of $\mathbb{F}^{\times}$. As a consequence, optimal unital $\mathbb{F}$-affine subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are classified, up to equivalence, by lists $([Q_{1}],\dots,[Q_{p}])$ of classes of nonisotropic Hermitian matrices over $\mathbb{D}$, under the relation of homo-star-congruence. In particular, if $\mathbb{D}\in\{\mathbb{C},\mathbb{H}\}$ and $\mathbb{F}=\mathbb{R}$, then we recover from the classification of Hermitian forms that every optimal unital $\mathbb{R}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is equivalent to $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ for a unique partition $n=n_{1}+\cdots+n_{p}$, and conversely $I_{n}+(\mathcal{S}\mathcal{H}_{n_{1}}(\mathbb{D})\vee\cdots\vee\mathcal{S}\mathcal{H}_{n_{p}}(\mathbb{D}))$ is an optimal unital $\mathbb{R}$-affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ for every partition $n=n_{1}+\cdots+n_{p}$.